OpenStudy (anonymous):

find the derivative of y with respect to x, t or theta as appropriate. y= ln(x)/x^6

OpenStudy (displayerror):

\[y=(x^{-6})*\ln x\] Use the product rule (derivative with respect to x) \[y'=(x^{-6})*(\frac{1}{x})+(\ln x)*(-6x^{-7})\]

OpenStudy (anonymous):

the answer is 1-7 ln(x)/ x^7 how can u get that?

OpenStudy (displayerror):

Simplify from what I typed and you should get the exact same answer. Remember that \[x^{-a} = \frac{1}{x^a}\]

OpenStudy (anonymous):

thank you